40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 19     Entry time: Fri Oct 26 17:34:43 2007
Author: waldman 
Type: Other 
Category: OMC 
Subject: OMC + earthquake stops 

[Chub, chris, Pinkesh, Sam]

Last night we hugn the OMC for the first time and came up with a bunch of pictures and some problems. Today we address some of the problems and, of course, make new problems. We replaced the flat slotted disks with the fitted slotted disks that are made to fit into the counterbore of the breadboard. This changed the balance slightly and required a more symmetric distribution of mass. It probably did not change the total mass very much. We did find that the amount of cable hanging down strongly affected the breadboard balance and may also have contributed to the changing balance.

We also attached earthquake stops and ran into a few problems:

  • The bottom plate of the EQ stops is too thick so that it bumps into the tombstones
  • The vertical member on the "waist" EQ stops is too close to the breadboard, possibly interfering with the REFL beam
  • The "waist" EQ stops are made from a thin plate that doesn't have enough thickness to mount helicoils in
  • Helicoil weren't loaded in the correct bottom EQ stops
  • The DCPD cable loops over the end EQ stop looking nasty but not actually making contact

However, with a little bit of jimmying, the EQ stops are arrayed at all points within a few mm of the breadboard. Meanwhile, Chub has cabled up all the satellite modules and DCPD modules and Pinkesh is working on getting data into the digital system so we can start playing games. Tonight, I intend to mount a laser in Rana's lab and fiber couple a beam into the 056 room so we can start testing the suspended OMC.
ELOG V3.1.3-